A beaker contains 7.85 g of Cu2O. A second beaker contains some FeCl3. If the two beakers contain the same numbers of formula units, what is the mass of the FeCl3?
1
Expert's answer
2021-10-05T01:31:42-0400

Solution:

Mass of Cu2O is 7.85 g

The molar mass of Cu2O is 143.09 g/mol

Therefore,

(7.85 g Cu2O) × (1 mol Cu2O / 143.09 g Cu2O) = 0.05486 mol Cu2O

n(Cu2O) = 0.05486 mol


Since the two beakers contain the same numbers of formula units, then:

N(Cu2O) = N(FeCl3)

n(Cu2O) × Na = n(FeCl3) × Na

where Na is Avogadro's number (6.022×1023 formula units mol-1)

Therefore,

n(Cu2O) = n(FeCl3)


n(FeCl3) = n(Cu2O) = 0.05486 mol

The molar mass of FeCl3 is 162.2 g/mol

Therefore,

(0.05486 mol FeCl3) × (162.2 g FeCl3 / 1 mol FeCl3) = 8.8983 g FeCl3 = 8.90 g FeCl3

Mass of FeCl3 is 8.90 g


Answer: The mass of FeCl3 is 8.90 grams.